Culver Corporation purchases a patent from Sunland Company on January 1, 2020, for $78,000. The patent has a remaining legal life of 16 years. Culver feels the patent will be useful for 10 years. Assume that at January 1, 2022, the carrying amount of the patent on Culver’s books is $62,400. In January, Culver spends $25,600 successfully defending a patent suit. Culver still feels the patent will be useful until the end of 2029.

Prepare the journal entries to record the $25,600 expenditure and 2022 amortization. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title to record expenditure of patents

enter a debit amount

enter a credit amount

enter an account title to record expenditure of patents

enter a debit amount

enter a credit amount

(To record expenditure of patents)

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

(To record amortization expense)
Answer #1
Date Account Titles and Explanation Debit Credit
January 1, 2022 Patent $25,600
Cash $25,600
(To record expenditure of patents)
December 31, 2022 Amortization expense - Patent $11,000
Accumulated amortization - Patent $11,000
(To record amortization expense)

Annual amortization of patent = Cost of patent /Useful life

= 78,000/10

= $7,800

$25,600 spent on successfully defending the patent suit will be amortized over the remaining useful life of patent.

Annual amortization of $25,600 = 25,600/8

= $3,200

Annual amortization of patent from 2022 onwards = 7,800+3,200

= $11,000

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks